OpenStudy (anonymous):

PPPLLLLEAASEE HEELLPPP!! Question 2 (Multiple Choice Worth 2 points) Solve 8x – 2 = 3 x ≈ 2.52 x ≈ 0.26 x ≈ 2.74 x ≈ 0.40 Question 7 (Multiple Choice Worth 2 points) Solve 5x + 5 = 9x x ≈ –3.63 x ≈ –0.28 x ≈ 0.07 x ≈ 13.69 Question 8 (Multiple Choice Worth 2 points) Solve 6x = 24 x ≈ 30.84 x ≈ 1.77 x ≈ 0.56 x ≈ 0.23 Question 11 (Multiple Choice Worth 2 points) Solve 7.74x + 3 = 10.2 x ≈ –2.15 x ≈ –0.47 x ≈ 0.03 x ≈ 0.28 Question 15 (Multiple Choice Worth 2 points) Solve 3x

OpenStudy (anonymous):

8x-2 = 3 8x-2+2 = 3+2 8x =5 x = 5/8 x=0.625 meaning there's now right answer in your choices, now do the same for the rest

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what exactly does the questions ask coz in all of them, the multiple choices are no where near the right answer!!!

OpenStudy (anonymous):

thees pics are the questions but pics of the questions! :P

OpenStudy (anonymous):

right we dealing with exponents, are you familiar with logs?

OpenStudy (anonymous):

Make sure when theres and exponent, put the ^ sign. Your first question should be 8^(x-2)=3.

OpenStudy (anonymous):

I'll do the first one then see if you can follow suite: \[8^{x-2} =3\] taking log of both sides: (x-2)log8 = log3 x-2 = log3/log8 x-2 = 0.528 x= 2.528 do you follow? remember the log is to base 10

OpenStudy (anonymous):

@cassie4math ????